OpenStudy (anonymous):

Related Time Rates Problem. AB & AC are two interacting roads enclosing an angle of measure 60 degrees. A bicycle moved on the road AB starting from A with uniform velocity of 16 Km/hr. 15 minutes later another bicycle moved on the road AC starting form A with uniform velocity 20 Km/hr. Find the rate of change of the distance between the two bicycles 15 minutes after starting of the second bicycle.

OpenStudy (amistre64):

the derivative of the pythag thrm should be useful

OpenStudy (amistre64):

 or maybe not :)

OpenStudy (amistre64):

law of cosines is a generalization pythag thrm and could be useful

OpenStudy (amistre64):

a^2 = b^2+c^2-2bc cos(A) 2a a' = 2b b' + 2c c' - 2b'c cos(A) -2bc' cos(A) a a' = b b' + c c' - b'c cos(A) - bc' cos(A) a' = (b b' + c c' - b'c cos(A) - bc' cos(A)) / a a' = (30(16)(16) + 15(20)(20) - (16)15(20) cos(30) - 30(16)(20) cos(30)) ----------------------------------------------------------- /(distance between the bikes)

OpenStudy (amistre64):

and distance between the bikes is just the solution of the law of cosines a^2 = b^2+c^2-2bc cos(A) a = sqrt(b^2+c^2-2bc cos(A)) a = sqrt(30^2 16^2 + 15^2 20^2 - 2(30)(15)(15)(20) cos(30))

OpenStudy (anonymous):

Thanks for your help. I appreciate it. :)